Please need your help in Solving the following system of linear equations
3x+2y=7
4x+3y=10

Answer by CubeyThePenguin(3113)

There are many ways to solve a system of equations, but we will use elimination.

3x + 2y = 7 (eqn 1)
4x + 3y = 10 (eqn 2)

Multiply eqn 1 by 3 and eqn 2 by 2:

9x + 6y = 21
8x + 6y = 20

Subtract to get x = 1 and y = 2.